A new cross-country water pipeline needs to be built at an estimated first cost of R 500 million. The consortium of cooperating companies has not fully decided the financial arrangements of this adventurous project. The WACC for similar projects has averaged 10% per year. Two financing options have been identified. The first requires an investment of 60% equity funds at 12% and a loan for the balance at an interest rate of 8% per year. The second option requires only 20% equity funds and the balance obtained by a massive international loan estimated to carry an interest rate of 13% per year. You are requested to decide which financing plan will result in the smaller average cost of capital?
Answer #1

Ans :

A new cross country water pipeline needs to be built at an estimated first cost = 500 million

The WACC for similar project has averaged = 10% per year

option 1 : 60% Equity funds at 12% per year

40% debt fund at 8% interest per year

option 2 : 20% Equity funds at 12% per year (similar )

80% debt funds at 13% interest per year

WACC = Cost of Equity * % Equity   + Cost of Debt * % Debt *(1-tax rate)

Here we assume tax rate is nil .

Option 1 : WACC =( 0.60*0.12 * 500) * 60% +   (0.40*0.08*500)* 40%

                                = 36 * 0.60 + 16*0.40

                                  = 28 million   Ans.

                              Or

                                 = (28/500) * 100 = 5.6 % Ans.

Option 2 : WACC =( 0.20*0.12 * 500) * 20% +   (0.80*0.13*500)* 80%

                                = 12 * 0.20 + 52*0.80

                                  = 44 million   Ans.

                              Or

                                 = (44/500) * 100 = 8.8 % Ans.

This ans. Applicable for low tax rate if tax rate is nil the both option is better then average similar project WACC but option 1 is gives low cost so that selected.

If tax rate is high then 45% (approx..) then as calculation option 2 is better then 1